OpenStudy (love.kat123):

Find the slope of the straight line that passes through (–2, –4) and (3, –5). A. m= -1/5 B. m = –1 C. m= 1/5 D. m = 1

OpenStudy (love.kat123):

@deepika.comet

OpenStudy (deepika.comet):

1 min

OpenStudy (love.kat123):

ok

OpenStudy (mokeira):



OpenStudy (mokeira):

hope you understood what i have done

OpenStudy (love.kat123):

my computer dont get pictures can u please explain @Mokeira

OpenStudy (mokeira):

\[m=\frac{ \Delta y }{\Delta x }\] \[m=\frac{ (-4--5) }{ -2-3 }\] \[m=\frac{ 1 }{-5 }\] Is that better

OpenStudy (love.kat123):

yes

OpenStudy (texaschic101):

slope formula : slope = (y2 - y1) / (x2 - x1) (-2,-4)...x1 = -2 and y1 = -4 (3,-5)....x2 = 3 and y2 = -5 now sub slope = (-5 - (-4) / (3 - (-2) slope = (-5 + 4) / (3 + 2) slope = -1/5 yep....its -1/5

OpenStudy (mokeira):

ok...hope you understood my steps

OpenStudy (love.kat123):

thank u

OpenStudy (texaschic101):

:) Thought it would help if you knew the slope formula rather then just writing numbers

OpenStudy (mokeira):

@texaschic101 true lol

OpenStudy (deepika.comet):

they explained it really good than me @Love.Kat123 got it now.!!